OpenStudy (abbycross167):

Can someone please help me with this math question? Please? I will fan and give a medal!!

OpenStudy (abbycross167):

Please help me to simplify this rational expression - (6x) (2y+2) ------ * ------- (y+1) (3)

OpenStudy (abbycross167):

@mathstudent55 do you think you could try to help me? please?

OpenStudy (wilsondanielle):

start by multiplying the fractions, and getting the product. Do you know how to multiply fractions?

jhonyy9 (jhonyy9):

why dont believ in my help ?

OpenStudy (abbycross167):

@wilsondanielle would I cross multiply? Like (6x)*(3)?

OpenStudy (wilsondanielle):

you cross when you divide - multiplying fractions is really easy. Multiple the top and then multiply the bottom! eg : 3x/2 * 4/2x = 12x/4x (=3)

OpenStudy (abbycross167):

Oh ok but where did you get the 4?

OpenStudy (wilsondanielle):

2*2x = 4. That's an example fraction - it has nothing to do with your actual fraction. What would the product of YOUR fraction be?

OpenStudy (abbycross167):

Oh! I was a little confused for a second!

OpenStudy (wilsondanielle):

no worries! once you have your product we can start simplifying, so just post it when you have it.

OpenStudy (abbycross167):

Would I do 6*2? For the top? The (2y+2) I don't understand

OpenStudy (wilsondanielle):

don't worry about splitting that up yet. you should wind up with something like this : 6x (2y+2) _________ 3 (y+1)

OpenStudy (wilsondanielle):

does that look familiar at all?

OpenStudy (abbycross167):

Somewhat, so it would be, 6x (2y+2) 12y ________ = ------------- 3 (y+1) 3 ?

OpenStudy (wilsondanielle):

close! it should be x, not y, and 12 can be divided by 3.

OpenStudy (abbycross167):

Oh ok so, (2y+2) 12x ________ = ------------- 3 (y+1) 3 12/3=4?

OpenStudy (wilsondanielle):

your final answer should still include an x - nothing is canceling it out.

OpenStudy (abbycross167):

Oh ok

OpenStudy (abbycross167):

So then what do we need to do?

OpenStudy (wilsondanielle):

just try the final simplification once more - you almost had it. What is 12x/3?

OpenStudy (abbycross167):

12x/3=3x

OpenStudy (wilsondanielle):

double check 12/3. It isn't 3...

OpenStudy (abbycross167):

4x sorry I meant to put that! Thank you for correcting me!

OpenStudy (wilsondanielle):

yep that's perfect! good job

OpenStudy (abbycross167):

Awesome!! thank you fr being a great teacher!! you really made it easy to understand!

OpenStudy (abbycross167):

I have two more that I am having some truble with would you mind helping me with them?

OpenStudy (wilsondanielle):

haha thank you! I'm glad you got it. Now you can even teach someone else!

OpenStudy (wilsondanielle):

I can help you with them. You can make new questions, message me, or post them here. Just tag me if you post them elsewhere!

OpenStudy (abbycross167):

yes sir/ ma'am, I'll post as a new question so you can get another medal